# Jean DELHAYE

### v1

@@ 3.3. Absolute continuity
In this paper, to compute limit profiles, we only consider the total variation distance. Just like in the classical setting, the total variation distance for states is much easier to compute when one of the state is absolutely continuous w.r.t the other. A state $\psi$ is said to be absolutely continuous w.r.t the Haar state if there exists $f\in\mathrm{L}^{1}(U_{N}^{+})$ such that $\psi(x)=h(fx)$ for all $x\in\mathcal{O}(U_{N}^{+})$. In this case, the total variation distance is easily expressed (see [16, Lem 2.6])
@@
Using the estimates from [13, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have
$\|f_{t}\|_{2}^{2}\leq\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}^{2}e^{-2t(\alpha-\beta/N)\lambda_{\mathbf{n}}}\leq 2\sum_{m\geq 1}\sum_{\begin{subarray}{c}\mathbf{n}=(n_{1},...,n_{p}),\\ n_{1}+...+n_{p}=m\end{subarray}}N^{m}e^{-2t(\alpha-\beta/N)m/N}\leq\sum_{m\geq 1}e^{-cm}<\infty.$
@@ 4. Restricting the study to a smaller algebra
$\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})\underset{N\to\infty}{\longrightarrow}e^{-c^{\prime}|\mathbf{n}|},\quad\hbox{with }c^{\prime}=c+\ln\sqrt{2}$
@@
This follows from the fact that $d_{n_{j}}e^{-t_{c}\lambda_{n_{j}}}\to e^{-\widetilde{c}n_{j}}$ as $N\to\infty$ for any $j\in\{1,...,p\}$ (see the proof of [17, Prop 3.10]). ∎
@@
The following proposition shows that for any $m$, the element $x_{m}$ is a linear combination of irreducible characters whose tuples are compositions of $m$. This implies not only that all the information of the limit-process is contained withing this algebra but it also gives the existence of a conditional expectation $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{00}$ leaving both the Haar state and the limit-process invariant (seeing as the $x_{m}$’s constitue an orthonormal family for the Haar state).
@@
Finally, let us end this subsection by mentioning that the closure of $\mathcal{O}(U_{N}^{+})_{00}$ in $C(U_{N}^{+})$ is isomorphic to $C([-N,N])$. Indeed, it suffices to see that the spectrum of $x/\sqrt{2}$ is $[-N,N]$. The inclusion $\sigma(x/\sqrt{2})\subset[-N,N]$ follows from the fact that $\|x/\sqrt{2}\|_{C(U_{N}^{+})}\leq N$. On the other hand, there exists a surjective C*-homomorphism $C(U_{N}^{+})\to C(O_{N}^{+}),u_{ij}\mapsto o_{ij}$. In particular, $x/\sqrt{2}$ is sent to $\chi_{1}$ which is well-known to have spectrum $[-N,N]$ (see [7, Lem 4.2]), hence the other inclusion.
@@ 5. The limit profile
For any $c>0$, we have
$d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),$
where $t_{c}=N\ln(\sqrt{2}N)/\alpha+cN/\alpha$.
@@
The proof of this result closely follows the reasoning in the proof [17, Prop 3.10], with a few modifications. Let us outline the elements that differ.
@@
Let $c>0$. It follows from Proposition 3.8 that $\widetilde{\psi}_{t_{c}}$ has an $\mathrm{L}^{1}$-density (at least asymptotically defined) given by
@@
In parallel, note that for any tuple $\mathbf{n}=(n_{1},...,n_{p})$, we have $d_{\mathbf{n}}e^{-t_{c}\lambda_{\mathbf{n}}}\to e^{-\widetilde{c}|\mathbf{n}|}$ as $N\to\infty$ where $\widetilde{c}=c+\ln\sqrt{2}$. This follows from the fact that $d_{n_{j}}e^{-t_{c}\lambda_{n_{j}}}\to e^{-\widetilde{c}n_{j}}$ as $N\to\infty$ for any $j\in\{1,...,p\}$ (see the proof of [17, Prop 3.10]). This further implies that for any given $m\in\mathbb{N}$, we have
@@
For any $c>0$, we have
$d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),$
where $t_{c}=N\ln(\sqrt{2}N)/\alpha+cN/\alpha$.
@@
It suffices to see that $d_{\mathrm{TV}}\big{(}\psi_{t_{c}},\widetilde{\psi}_{t_{c}}\big{)}\to 0$ as $N\to\infty$. This is obvious as
@@
where the last inequality holds for $N$ large enough (similarly to the proof of Proposition 3.8). By dominated convergence, we may exchange summation and limit in $N$ yielding the convergence. ∎
@@ 6. Non-uniqueness of limit profiles
$\hat{\eta}_{0}(P_{n})=\int_{-2}^{2}\frac{2P_{n}(x)}{2-x^{2}}\mathrm{d}\nu_{\mathrm{SC}}(x)=\lim_{c\to 0}\hat{\eta}_{c}(P_{n})=\left\{\begin{array}[]{cl}1&\hbox{if }n\hbox{ is even}\\ 0&\hbox{else},\end{array}\right.\quad n\in\mathbb{N}.$
@@
Now to prove the limit profile recall that we can see $\varphi_{t_{N}}$ and $h$ as their measure counterparts $\mu_{t_{N}}$ and $\nu_{\mathrm{SC}}$ via the isomorphism $\chi_{n}\mapsto P_{n}$. It remains to see that the absolutely continuous part of $\mu_{t_{N}}$ w.r.t $\nu_{\mathrm{SC}}$ tends to that of $\hat{\eta}_{c}$ as $N\to\infty$ in total variation distance which can be done in a similar fashion as in the proof of [17, Thm 3.9]. ∎
Note that the functions $c\mapsto d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}})$ and $c\mapsto d_{\mathrm{TV}}(\hat{\eta}_{c},\nu_{\mathrm{SC}})$ are not equivalent modulo affine transformation. It follows from Remark 2.8 that the profile we exhibited is different.
@@
In this appendix, we give a description of Gaussian centralized generating functionals on $O_{N}^{+}$ and $U_{N}^{+}$. We will therefore focus our study on the latter quantum groups, we refer the reader to [14, Sec 1.5] for definitions in greater generality.
@@
[12, Prop 2.7] Let $L$ be a Gaussian generating functional on $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$ Then, for any $a_{1},...,a_{n}\in\mathcal{O}(\mathbb{G})$ we have
@@
Let us first note that any Gaussian cocycle $\eta$ defined on a compact quantum group satisfies the relation $\eta\circ S+\eta=0$ (see the beginning of the proof of [12, Thm 3.11]).

### v2

@@ 3.3. Absolute continuity
In this paper, to compute limit profiles, we only consider the total variation distance. Just like in the classical setting, the total variation distance for states is much easier to compute when one of the state is absolutely continuous w.r.t the other. A state $\psi$ is said to be absolutely continuous w.r.t the Haar state if there exists $f\in\mathrm{L}^{1}(U_{N}^{+})$ such that $\psi(x)=h(fx)$ for all $x\in\mathcal{O}(U_{N}^{+})$. In this case, the total variation distance is easily expressed (see [21, Lem 2.6])
@@
Using the estimates from [17, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have
$\|f_{t}\|_{2}^{2}\leq\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}^{2}e^{-2t(\alpha-\beta/N)\lambda_{\mathbf{n}}}\leq 2\sum_{m\geq 1}\sum_{\begin{subarray}{c}\mathbf{n}=(n_{1},...,n_{p}),\\ n_{1}+...+n_{p}=m\end{subarray}}N^{2m}e^{-2t(\alpha-\beta/N)m/N}\leq\sum_{m\geq 1}e^{-cm}<\infty.$
@@ 4. Restricting the study to a smaller algebra
$\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})\underset{N\to\infty}{\longrightarrow}e^{-\widetilde{c}|\mathbf{n}|},\quad\hbox{with }\widetilde{c}=c+\ln\sqrt{2}$
@@
This follows from the fact that $d_{n_{j}}e^{-t_{c}\lambda_{n_{j}}}\to e^{-\widetilde{c}n_{j}}$ as $N\to\infty$ for any $j\in\{1,...,p\}$ (see the proof of [23, Prop 3.10]). ∎
@@
The following proposition shows that for any $m$, the element $x_{m}$ is a linear combination of irreducible characters whose tuples are compositions of $m$. This implies not only that all the information of the limit-process is contained within this algebra but it also gives the existence of a conditional expectation $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{00}$ leaving both the Haar state and the limit-process invariant (seeing as the $x_{m}$’s constitute an orthonormal family for the Haar state).
@@
Finally, let us end this subsection by mentioning that the closure of $\mathcal{O}(U_{N}^{+})_{00}$ in $C(U_{N}^{+})$ is isomorphic to $C([-N,N])$. Indeed, it suffices to see that the spectrum of $x/\sqrt{2}$ is $[-N,N]$. The inclusion $\sigma(x/\sqrt{2})\subset[-N,N]$ follows from the fact that $\|x/\sqrt{2}\|_{C(U_{N}^{+})}\leq N$. On the other hand, there exists a surjective C*-homomorphism $C(U_{N}^{+})\to C(O_{N}^{+}),u_{ij}\mapsto o_{ij}$. In particular, $x/\sqrt{2}$ is sent to $\chi_{1}$ which is well-known to have spectrum $[-N,N]$ (see [9, Lem 4.2]), hence the other inclusion.
@@ 5. The limit profile
We have
$d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\geq d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c<0,$
@@
The proof of the first convergence closely follows the reasoning in the proof [23, Prop 3.10], with a few modifications. Let us outline the elements that differ. Let $c>0$. It follows from Proposition 3.8 that $\widetilde{\psi}_{t_{c}}$ has an $\mathrm{L}^{1}$-density (at least asymptotically defined) given by
@@
In parallel, note that for any tuple $\mathbf{n}=(n_{1},...,n_{p})$, we have $d_{\mathbf{n}}e^{-t_{c}\lambda_{\mathbf{n}}}\to e^{-\widetilde{c}|\mathbf{n}|}$ as $N\to\infty$ where $\widetilde{c}=c+\ln\sqrt{2}$ (this follows from Lemma 4.1 and Proposition 4.3). This further implies that for any given $m\in\mathbb{N}$, we have
@@
Now, for the second result, fix $c<0$ and consider the states $\widetilde{\psi}_{t_{c}}$ and $h$ in their measure form as $\mu_{t_{c}}$ and $\nu_{\mathrm{SC}}$. We have already established that $\mu_{t_{c}}$ converges to $\eta_{c}$ in moments as $N\to\infty$. Taking $N$ large enough so that $t_{c}\beta\leq-cN^{2}$, using the estimates from [17, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have for any $m\in\mathbb{N}$
$\mu_{t_{c}}(P_{m})\leq 2^{m/2}N^{m}e^{-t_{c}(1-\beta/N)m/N}\leq e^{-2cm}=\eta_{2c}(P_{m}).$
In particular, this implies that:
$\mu_{t_{c}}(X^{2n})\leq\eta_{2c}(X^{2n}),\quad\text{for all }n\in\mathbb{N}.$
This follows from the fact that $X^{2n}$ can be expressed as a linear combination of $P_{m}$’s with positive coefficients. Since $\eta_{2c}$ is supported on the interval $[-\gamma_{2c},\gamma_{2c}]$, where $\gamma_{2c}=e^{2c}+e^{-2c}$, it follows that $\mu_{t_{c}}$ must also be supported on this interval (as the support of a measure can be estimated by bounding its moments; see, for example, the beginning of [32, Lect 2]). Given that moment convergence with compact support implies weak convergence, we have:
$\mu_{t_{c}}(B)\xrightarrow[N\to\infty]{}\eta_{c}(B),$
for any continuity set $B$ of $\eta_{c}$, i.e., a set such that $\eta_{c}(\partial B)=0$. The final step is to identify such a set that maximizes the total variation distance between $\eta_{c}$ and $\nu_{\mathrm{SC}}$. It suffices to take:
$B_{c}=\{x\in[-2,2]:g_{c}(x)\leq 1\},$
where $g_{c}$ is the density of the absolutely continuous part of $\eta_{c}$ with respect to the semicircular distribution. With this choice of $B_{c}$, we now conclude the argument.
$d_{\mathrm{TV}}(\mu_{t_{c}},\nu_{\mathrm{SC}})\geq\nu_{\mathrm{SC}}(B_{c})-\mu_{t_{c}}(B_{c})\underset{N\to\infty}{\longrightarrow}\nu_{\mathrm{SC}}(B_{c})-\eta_{c}(B_{c})=d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}}).$
@@
We have
$d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\geq d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c<0,$
where $t_{c}=N\ln(\sqrt{2}N)/\alpha+cN/\alpha$. In particular, the Brownian motion has cutoff at time $t_{N}=N\ln(\sqrt{2}N)/\alpha$.
@@
Fix $c>0$, for the first convergence it suffices to see that $d_{\mathrm{TV}}\big{(}\psi_{t_{c}},\widetilde{\psi}_{t_{c}}\big{)}\to 0$ as $N\to\infty$. This is obvious as
@@
where the last inequality holds for $N$ large enough (similarly to the proof of Proposition 3.8). By dominated convergence, we may exchange summation and limit in $N$ yielding the convergence.
For the second one, simply note that for any $c\in\mathbb{R}$, we have
$d_{\mathrm{TV}}(\psi_{t_{c}},h)=\frac{1}{2}\|\psi_{t_{c}}-h\|_{FS}\geq\frac{1}{2}\|\mathbb{F}\circ\psi_{t_{c}}-\mathbb{F}\circ h\|_{FS}=d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h).$
Let us conclude this section by discussing the result. The limit profile on the right ($c>0$) is straightforward to compute due to the inherent absolute continuity in this region. However, outside of this region, absolute continuity is lost. It is worth recalling that such a loss of absolute continuity was already observed in [23], where all processes studied exhibited an atom as the singular part when absolute continuity was no longer present. P. Biane, in [6, Sec 12.2], introduced an interesting method for recovering the absolutely continuous part of a measure when only its moments are known, provided that one can identify its atom. In our case, the singular part is more complex than a single atom, and it remains unclear whether Biane’s method could be applied here. We conjecture, however, that the singular part lies outside the support of $\nu_{\mathrm{SC}}$, and that there is true $\mathrm{L}^{1}$-convergence. Specifically, we believe that even for $c<0$, the absolutely continuous part of the process converges to that of $\eta_{c}$ in $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$.
@@ 6. Non-uniqueness of limit profiles
$\hat{\eta}_{0}(P_{n})=\int_{-2}^{2}\frac{2P_{n}(x)}{2-x^{2}}\mathrm{d}\nu_{\mathrm{SC}}(x)=\lim_{c\to 0^{+}}\hat{\eta}_{c}(P_{n})=\left\{\begin{array}[]{cl}1&\hbox{if }n\hbox{ is even}\\ 0&\hbox{else},\end{array}\right.\quad n\in\mathbb{N}.$
@@
Now to prove the limit profile recall that we can see $\varphi_{t_{N}}$ and $h$ as their measure counterparts $\mu_{t_{N}}$ and $\nu_{\mathrm{SC}}$ via the isomorphism $\chi_{n}\mapsto P_{n}$. It remains to see that the absolutely continuous part of $\mu_{t_{N}}$ w.r.t $\nu_{\mathrm{SC}}$ tends to that of $\hat{\eta}_{c}$ as $N\to\infty$ in total variation distance which can be done in a similar fashion as in the proof of [23, Thm 3.9]. ∎
Observe that the functions $f_{1}:c\mapsto d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}})$ and $f_{2}:c\mapsto d_{\mathrm{TV}}(\hat{\eta}_{c},\nu_{\mathrm{SC}})$ can be explicitly computed through integration, resulting in the following expressions:
$f_{1}(c)=\begin{cases}\frac{2(1-e^{2c})}{\pi}\arcsin\left(\frac{e^{-c}}{2}\right)+\frac{e^{-2c}+2}{4\pi}\sqrt{4e^{2c}-1},&\text{if }c\geq-\ln 2,\\ 1-e^{2c},&\text{if }c<-\ln 2,\end{cases}$
and
$f_{2}(c)=\begin{cases}\frac{\sqrt{1+e^{-2c}}\sqrt{3-e^{-2c}}}{2\pi}+\frac{e^{2c}-1}{\pi}\left(\arctan\left(\frac{\sqrt{1+e^{-2c}}(e^{c}-e^{-c})}{(e^{c}+e^{-c})\sqrt{3-e^{-2c}}}\right)-\arcsin\left(\frac{\sqrt{1+e^{-2c}}}{2}\right)\right),&\text{if }c\geq-\frac{\ln 3}{2},\\ 1-e^{2c},&\text{if }c<-\frac{\ln 3}{2}.\end{cases}$
Figure 1. Graph comparing the functions $f_{1}$ (red) and $f_{2}$ (blue).
They agree on infinitely many points but are indeed distinct. This implies that they are not equivalent modulo affine transformation. It follows from Remark 2.8 that the profile we exhibited is different.
Interestingly, although we might have anticipated that the new process would result in a ’faster shuffle’, in the sense that $f_{2}\leq f_{1}$, the two profiles initially overlap for a significant range before the impact of the ’faster shuffle’ becomes apparent.
@@
In this appendix, we give a description of Gaussian centralized generating functionals on $O_{N}^{+}$ and $U_{N}^{+}$. We will therefore focus our study on the latter quantum groups, we refer the reader to [19, Sec 1.5] for definitions in greater generality.
@@
[16, Prop 2.7] Let $L$ be a Gaussian generating functional on $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$ Then, for any $a_{1},...,a_{n}\in\mathcal{O}(\mathbb{G})$ we have
@@
Let us first note that any Gaussian cocycle $\eta$ defined on a compact quantum group satisfies the relation $\eta\circ S+\eta=0$ (see the beginning of the proof of [16, Thm 3.11]).